# Constraint Breeds Generalization: Temporal Dynamics as an Inductive Bias

### v1

@@ Abstract
Conventional deep learning prioritizes unconstrained optimization, yet biological systems operate under strict metabolic constraints. We propose that these physical constraints shape dynamics to function not as limitations, but as a temporal inductive bias that breeds generalization. Through a phase-space analysis of signal propagation, we reveal a fundamental asymmetry: expansive dynamics amplify noise, whereas proper dissipative dynamics compress phase space that aligns with the network’s spectral bias, compelling the abstraction of invariant features. This condition can be imposed externally via input encoding, or intrinsically through the network’s own temporal dynamics. Both pathways require architectures capable of temporal integration and proper constraints to decode induced invariants, whereas static architectures fail to capitalize on temporal structure. Through comprehensive evaluations across supervised classification, unsupervised reconstruction, and zero-shot reinforcement learning, we demonstrate that a critical "transition" regime maximizes generalization capability. These findings establish dynamical constraints as a distinct class of inductive bias, suggesting that robust AI development requires not only scaling and removing limitations, but computationally mastering the temporal characteristics that naturally promote generalization.
@@ 1 Introduction
Modern deep learning paradigms largely abstract the temporal dimension, simplify time to a spatial index to enable efficient parallelization[lecun2015deep, krizhevsky2012imagenet, hooker2021hardware]. This modification contrasts sharply with biological systems, where temporal dynamics are intrinsic to computation and operate under severe energetic constraints[buzsaki2006rhythms, shankar2025bridging]. While current AI equates generalization with unconstrained scaling[kaplan2020scaling, neyshabur2017exploring], the brain’s strict metabolic budget forces neural dynamics into a dissipative regime characterized by sparse, temporally precise activity[attwell2001energy, olshausen1996emergence, van2001rate]. We propose that this energy-efficiency trade-off functions not as a limitation, but also as a fundamental inductive bias: by naturally compressing the solution space over time, these physical constraints compel the system to abstract robust invariants from noisy sensory streams.
Specifically, we hypothesize that networks trained on temporal-encoded, well-structured inputs will inherently learn representations invariant to task-irrelevant variations, treating time not merely as a dimension for indexing but as a physical resource for computation. Building on recent work in dynamical alignment[chen2025dynamical], which demonstrated that SNN computational modes can be sculpted by input dynamics, we posit that temporal constraints serve as an active regularizer. To instantiate this principle computationally, we leverage Spiking Neural Networks (SNNs), which provide the necessary theoretical foundation for temporal processing and biological plausibility[roy2019towards, eshraghian2023training].
We designed experiments sequentially based on emergent observations and established that constraint, specifically structured dissipative temporal dynamics, breeds generalization across:
@@
    Behavioral level: In reinforcement learning, where inputs are inherently temporal, we validate the constraint principle through two complementary pathways: encoding-level constraints (consistent with Experiments 1 and 2) and architecture-level constraints (via neuronal leak mechanisms). Both pathways achieve superior zero-shot transfer to unseen physical environments, outperforming baseline architectures (Experiment 3).
Furthermore, we discover that this generalization capability requires the co-dependence of temporal input dynamics and processing architecture: conventional time-indexing architectures (which treat time merely as a dimension) fail to benefit from these dynamical constraints even when provided with identical inputs. Mechanistically, spectral analysis reveals that the proper contractive input trajectory (we called "transition regime") emerges as a "low-frequency, high-entropy" signature, which aligns with the theory of spectral bias in neural networks[rahaman2019spectral]. These findings suggest dynamical constraints as a distinct class of inductive bias[goyal2022inductive] working on the temporal dimension. Unlike explicit, hand-crafted spatial regularizers (e.g., data augmentation[hernandez2018data, balestriero2022effects] or architectural design[battaglia2018relational]), this implicit temporal approach operates through a physical phase-space contraction (defined by $\Sigma\lambda_{i}$) process that actively induces the network to learn stable, invariant features over time and consistently robust across representational, structural, and behavioral levels.
@@ 2.1 Method: Dynamical Constraints via Dual Pathways
We implement dissipative constraints through two complementary pathways, thereby establishing the framework’s generality. Both pathways control the same physical quantity: the rate of phase space contraction during the temporal evolution of signals, measured by the global Lyapunov sum ($\Sigma\lambda_{i}$)[wolf1985determining]. As established in prior work[chen2025dynamical], the computational capability of downstream networks is primarily governed by the global phase space contraction ($\Sigma\lambda_{i}$) of the input, rather than local chaotic divergence ($\lambda_{max}$).
Encoding-Level Constraints. For tasks with static inputs (Experiments 1-2), we transform features $\mathbf{x}\in\mathbb{R}^{d}$ into temporal trajectories using a parametrically controllable duffing oscillator system[kovacic2011duffing]. This design allows us to tune the global phase space dynamics via a single parameter $\delta$ within the same system to avoid geometric confounds inherent when comparing distinct chaotic systems. Each input dimension initializes a three-dimensional dynamical system that evolves for time $T$ with $N$ discrete steps, producing trajectories $\phi_{\delta}(\mathbf{x},t)\in\mathbb{R}^{d\times N\times 3}$. The parameter $\delta$ controls the system’s phase space behavior:
@@
Architecture-Level Constraints. For tasks with inherently temporal inputs (Experiment 3), we also leverage the SNN’s intrinsic dissipation mechanism of Leaky Integrate-and-Fire (LIF) neurons[eshraghian2023training]. The membrane leak parameter $\beta$ governs temporal integration within the network:
@@
Specifically, $\beta$ governs the effective integration window ($\tau_{\text{mem}}=-1/\ln(\beta)$), acting as the architectural counterpart to the encoding parameter $\delta$. High $\beta$ ($\approx 1.0$) preserves information across timesteps (weak constraint), while low $\beta$ ($\approx 0.1$) rapidly dissipates internal states (strong constraint). This setup allows Experiment 3 to validate that our findings derive from the constraint principle itself, rather than properties specific to the input encoding.
2.2 Experiment 1: Emergent Generalization in Classification
@@
We first investigate representational transfer[yosinski2014transferable, deng2022strong] as a direct proxy for invariance. The rationale is straightforward: If dissipative constraints indeed lead to the abstraction of robust features, a network trained on one dynamical encoding should generalize to others, analogous to domain adaptation but operating within the spectrum of dynamical regimes.
@@
Specifically, we test whether networks trained under different dynamical encodings exhibit generalization across regimes. We designed a cross-encoding evaluation protocol evaluating five networks on the handwritten digits dataset (64 features, 10 classes)[optical_recognition_of_handwritten_digits_80]: (1) SNNs with LIF Neuron, compared against recurrent controls: (2) Long Short-Term Memory networks (LSTMs)[hochreiter1997long], (3) Recurrent Neural Networks (RNNs); as well as static baselines: (4) Multi-Layer Perceptrons (MLPs) processing only the final timestep (Last-T), (5) MLPs processing time-averaged inputs (Avg-Pool). All networks share an identical three-layer fully-connected topology to isolate algorithmic differences. We constructed a $12\times 12$ generalization matrix by training each architecture on a specific dynamical regime (governed by $\delta_{\text{train}}$) and evaluating it across the full spectrum of test dynamics ($\delta_{\text{test}}\in\{-1.5,\dots,10.0\}$). Detailed experiment settings are provided in Appendix A.1.
@@
Figure 1: Cross-encoding generalization with hierarchical variance reduction. (a-e) 3D generalization landscapes (mean accuracy with deviation, $n=10$). (c) SNNs exhibit a robust generalization in the transition regime ($\delta\approx 0-2.0$) across the spectrum. This contrasts with the diagonal-only ridges seen in expansive regimes ($\delta<0$) and in baseline architectures (Last-T MLP, LSTM, RNN; b, d, e), and Avg-Pool MLP (a) shows partial, unstable generalization (see Appendix Tables 69). (f) Out-of-Distribution accuracy strongly anti-correlates with Layer 1 neural variability (CV; r=0.962r=-0.962). Error bars: $\pm 1$ SEM. (g) Hierarchical variance reduction (L1 $\to$ L3) emerges in SNN across regimes.
@@
Asymmetric Generalization Landscapes. Figure 1 presents the resulting generalization capability as a 3D topographic landscape. We define In-Distribution (ID) accuracy as the performance on the training encoding (the diagonal ridges where $\delta_{\text{train}}\approx\delta_{\text{test}}$), and Out-of-Distribution (OOD) accuracy as the performance across the rest of the spectrum. While all architectures maintain high ID performance (indicated by the peaks of the ridges), SNNs (Figure 1c) reveal a unique, asymmetric generalization capability: Networks trained in expansive regimes ($\delta_{\text{train}}<0$) form sharp, narrow ridges. They achieve high ID accuracy but collapse immediately in the $\delta_{\text{test}}>0$ region. This behavior is characteristic of memorization without transfer. In contrast, SNNs trained in the transition ($\delta_{\text{train}}\approx 0$) and dissipative ($\delta_{\text{train}}>2$) regimes form a robust "generalization plateau." These ridges maintain high altitude across the entire $\delta_{\text{test}}$ axis. Notably, the transition zone ($\delta_{\text{train}}\approx\text{0-2}$) achieves the strongest overall generalization, effectively decoding inputs from all other dynamical regimes.
This plateau is absent in recurrent controls (LSTMs, RNNs) and stateless MLPs (Last-T), which show OOD failure (collapsed ridges) when tested away from their training diagonal. The Avg-Pool MLP shows partial generalization but with significantly higher volatility (see Appendix Tables 59). This suggests that the mechanistic shortcut of averaging over the temporal dimension diminishes generalization capacity, as it primarily arises from a particular temporal structure of trajectories decoded by SNN’s integral dynamics, rather than from mere temporal aggregation.
@@
Emergent Hierarchical Variance Reduction. How do SNNs translate these dynamical constraints into robust representations? We hypothesized that dissipative constraints drive the network to suppress noise and amplify stable features. To quantify this, we measured the neural response stability via the coefficient of variation ($CV=\sigma/\mu$) of mean firing rates across test encodings. Figure 1f reveals a mechanistic insight: neural response variability (CV) strongly anti-correlates with OOD accuracy across the dynamical spectrum (r=0.96r=-0.96; see Appendix Table 10).
@@
Furthermore, this stabilization is not uniform but forms hierarchically over layers. As shown in Figure 1g, SNNs exhibit a monotonic reduction in variability across depth (L1 $\to$ L3). The representation stability in the deepest layer (Layer 3) shows the starkest contrast: a high fluctuating variance observed under the expansion regime ($0.470\pm 0.288$, see Appendix Table 11) suggests that the network memorizes specific dynamical trajectories. In contrast, dissipative constraints drive this final representation toward a tightly converged state ($0.307\pm 0.051$), indicating the formation of invariant representation (Appendix Table 11). This emergent phenomenon is highly similar to the processing hierarchy found in both standard deep networks[yosinski2014transferable] and the biological vision system (primate ventral stream), where raw, stimulus-specific volatility (akin to V1 responses) is progressively distilled into abstract, invariant representations (as in the inferotemporal cortex)[dicarlo2012does, akella2025deciphering]. However, unlike previous works that primarily attribute this invariance to prior spatial structure designs (e.g., pooling or strided convolutions[yamins2014performance]), our results demonstrate that such hierarchical robustness can arise purely from intrinsic temporal dissipation.
2.3 Experiment 2: Spontaneous Induction of Structured Features
@@
Previous observation from the experiment motivates us for subsequent inquiry: Does this variance reduction spontaneously induce organized feature representations? While classical theories posit sparse coding as an explicit optimization objective for cortical receptive fields [olshausen1996emergence, barlow1961possible], we hypothesize that dissipative dynamics act as a fundamental constraint that naturally induces the emergence of invariant, structured features. In other words, representational structure is not the result of a designed objective but an emergent property of phase space compression.
To test this, we trained a SNN autoencoder to reconstruct static image patches from temporally encoded inputs. The architecture features a LIF bottleneck layer between a linear encoder ($W_{\text{enc}}$) and decoder ($W_{\text{dec}}$). Critically, the decoder reconstructs from the temporal sum of spikes: $\mathbf{x}_{\text{recon}}=W_{\text{dec}}\sum_{t=1}^{T}S(t)$. We followed standard comparative coding schemes[auge2021survey, guo2021neural] and compared eight temporal encoding strategies on CIFAR-10 patches[krizhevsky2009learning], including Baseline (static), Random (temporal jitter), Linear (rate coding), Poisson (stochastic spiking), and four dynamical regimes (Expansive, Critical, Transition, Dissipative). The network minimized a composite loss balancing reconstruction fidelity and code sparsity. We quantified the organization of learned spatial features at the bottleneck using RF Standard Deviation ($\sigma_{\text{RF}}$) (distinguishing high-contrast structured filters from low-variance noise), alongside reconstruction and sparsity losses. Finally, we validated robustness through a three-fold evaluation: (1) varying the sparsity weight $\lambda$ (specifically including the unconstrained $\lambda=0$ case); (2) scaling experiments to high-resolution STL-10 images[coates2011analysis]; and (3) verifying dynamical universality across alternative physical systems (Lorenz[lorenz2017deterministic] and Thomas attractors[thomas1999deterministic]). Detailed experiment protocols are provided in Appendix A.2.
@@
Structure Emergence under Dynamical Constraints Figure 2 provides a qualitative visualization of feature emergence. Transition regime ($\delta\approx 2.0$) spontaneously produce receptive fields with clear spatial organization: localized, oriented filters resembling the biological orientation maps of the visual cortex. In contrast, baseline, random, and expansive conditions yield homogeneous, unstructured weight distributions resembling salt-and-pepper noise. Table 1 quantifies these distinct organizations across three metrics:
@@
  • Expansive dynamics ($\delta=-1.5$): Achieves excellent reconstruction ($0.0013$) but fails catastrophically at sparsity ($1.72$), which means that dynamics amplify temporal variations, allowing the network to "memorize" inputs via high-rate firing without learning compressive features.
@@
  • Poisson encoding: Demonstrates the opposite extreme with near-perfect sparsity ($10^{-5}$) but minimal structure ($0.049$). This proves that sparse firing alone is insufficient; temporal stability is required to induce structural differentiation.
@@
  • Transition dynamics ($\delta\approx 2.0$): Presents an optimal regime. Its contractive phase space simultaneously reduces activation volume (sparsity $0.0009$), suppresses noise to enable clean filters, and amplifies stable features (structure $0.206$, $\mathbf{4.9\times}$ higher than alternatives).
@@
Further analysis confirms that even without sparsity penalty ($\lambda=0$, see Appendix A.2), the Transition regime still spontaneously induces structured representations ($\sigma_{\text{RF}}\approx 0.177$), whereas others collapse into noise. In this regime, structure emerges intrinsically by the dynamical constraint, independent of explicit regularization. This offers a principled hypothesis for biological sparse coding: under dissipative metabolic constraints, structured representations emerge not as a pre-programmed objective, but as a canonical solution to learning under temporal constraints[attwell2001energy].
@@
Mechanism This structural emergence raises a fundamental question: What specific physical property of the transition dynamics drives this outcome? Spectral analysis confirms that the transition regime ($\delta\approx 2$) occupies a unique computational niche defined by a "High-Entropy, Low-Frequency" signature (Fig. 5, see Appendix 4). This signature is scale-invariant (Fig. 5B), ensuring that the network receives information-rich topological structures within a slow-varying envelope. By physically restricting input trajectories to a low spectral centroid, transition dynamics function as an intrinsic spectral shaper that aligns input complexity with the spectral bias of neural networks[rahaman2019spectral], which also explains the asymmetric generalization observed in previous experiments. This mechanism is robust: a high-resolution parameter sweep reveals that this structural emergence forms a robust ridge that persists (see Appendix 3). This confirms that feature organization arises from the transition dynamics itself, acting as a physical constraint that induces structure more naturally than hand-crafted than explicit regularization.
@@
Functionally, the observed structure (Figure 2) mirrors the spatial antagonism found in biological vision[olshausen1996emergence, ferster2000neural] and reflects a correlation-driven consolidation process: By suppressing high-frequency noise while preserving low-frequency structure, transition dynamics amplify spatiotemporal correlations. This selective amplification biases the optimization landscape toward grouping co-varying features, yielding outcomes that are much like Hebbian association[hebb2005organization, barlow1961possible] yet driven primarily by the temporal statistics of the input rather than explicit plasticity rules. Conversely, high-frequency chaos in the expansive regime decorrelates inputs and prevents such consolidation over time, while static or other baselines lack the necessary temporal continuity to induce structural binding. This suggests that biological dissipative constraints serve a computational role beyond energy efficiency: by naturally filtering temporal correlations, they provide a mechanistic basis for structured connectivity[buzsaki2006rhythms].
@@
Furthermore, this structural emergence corroborates recent work on diffusion dynamics, which suggests a critical gap between memorization and generative abstraction[bonnaire2025diffusion]. Our transition dynamics position the learning process on the generative side of this divide: while the expansive regime with high-frequency chaos allows the network to memorize transient noise for rapid error reduction (the "memorization phase"), the transition regime physically imposes a temporal information bottleneck (See Table 4 and Fig. 5 in Appendix) and implements Slow Feature Analysis[wiskott2002slow] through dynamics alone, which makes the system discard transient variations and rely on the underlying generative invariants. This re-interprets the "performance valley" observed in prior work[chen2025dynamical] not as a failure, but as the necessary cost of abstraction.
These lines of evidence unify our findings: structured dissipative dynamics function as an active ordering principle to enable both invariance (Exp 1) and organization (Exp 2). They are not distinct phenomena but different manifestations of constraint-driven induction, which naturally leads to the third test: if low-level invariance and structural emergence both stem from temporal constraint, does this translate to high-level behavioral robustness?

### v2

@@ Abstract
Modern deep learning prioritizes unconstrained scaling and optimization, yet biological systems exploit temporal dynamics under strict metabolic constraints. We propose that dissipative dynamics function as a unique class of inductive bias, distinct from spatial regularizers; temporal phase-space contraction compels networks to discard transient noise and consolidate invariant representations over time. Through phase-space analysis, we identify a critical regime that compels signal evolution to align with neural networks’ spectral bias, maximizing generalization across three levels: representational (cross-encoding classification), structural (spontaneous emergence of receptive fields), and behavioral (zero-shot transfer in reinforcement learning). Importantly, this capability requires temporal integration to decode invariants that static architectures perceive as noise. Our findings reframe temporal irreversibility from physical limitation to computational principle: robust AI development requires not merely scaling, but mastering the temporal characteristics that naturally promote generalization.
@@ 1 Introduction
Modern deep learning paradigms largely abstract away the temporal dimension, flattening time into a static spatial index to facilitate efficient parallelization [34, 36, 27]. This simplified view contrasts sharply with biological systems, where temporal dynamics are intrinsic to computation under severe energy budget [12, 52]. Unlike current AI paradigms that equate generalization with unconstrained scaling [42, 28], the brain’s strict metabolic demands force neural dynamics into a dissipative regime characterized by sparse, temporally precise activities [43, 3, 56]. Such energy-efficiency selection pressure serves not merely as a limitation, but as a fundamental organizing principle: by naturally compressing the solution space over time, this physical constraint compels the system to abstract robust representations from noisy sensory streams.
This observation motivates our central hypothesis: temporal structure, when properly constrained, functions as an active regularizer for generalization. Specifically, we propose that networks processing temporally encoded signals under dissipative dynamics will learn representations inherently invariant to task-irrelevant variations. To incorporate this principle computationally, we leverage Spiking Neural Networks (SNNs) as the computational substrate, as they provide the necessary theoretical foundation and biological plausibility [49, 19]. Through a sequence of experiments driven by emergent observations, we establish that structured dissipative constraints actively induce stable, invariant features that remain consistently robust across:
@@
    Behavioral level: In reinforcement learning, where inputs are inherently temporal, we validate the principle through two complementary pathways: encoding-level constraints (consistent with Experiments 1 and 2) and architecture-level constraints (via network internal state dissipation). Both pathways achieve superior zero-shot transfer to unseen physical environments, outperforming baseline architectures (Experiment 3).
Beyond these empirical findings, we discover that this generalization capability requires a co-dependence of signal dynamics and networks capable of temporal integration: conventional time-indexing architectures fail to benefit from this principle even when provided with identical inputs. Mechanistically, spectral analysis reveals that the proper contractive signal trajectory (termed "transition regime") emerges as a "low-frequency, high-entropy" signature, aligning with the spectral bias of neural networks [48]. These results suggest that dynamical constraints act as a distinct class of inductive bias [21] that exploits temporal irreversibility as a critical computational principle. Unlike explicit, hand-crafted spatial regularizers (e.g., data augmentation [25, 6] or architectural priors [8]), this implicit dissipative process operates through physical phase-space contraction (defined by $\Sigma\lambda_{i}$), suggesting that robust generalization may emerge not from removing constraints, but from computationally mastering them.
@@ 2.1 Method: Dynamical Constraints via Dual Pathways
To test whether this principle is implementation-agnostic, we realize dissipative constraints through two complementary pathways that both govern the same physical quantity: the global Lyapunov sum ($\Sigma\lambda_{i}$), which quantifies phase space contraction during signal evolution [60]. Prior work has established that downstream computational capability depends primarily on this global contraction rate rather than local chaotic divergence ($\lambda_{max}$[18, 13].
Encoding-Level Constraints. For tasks with static inputs (Experiments 1-2), we transform features $\mathbf{x}\in\mathbb{R}^{d}$ into temporal trajectories using a parametrically controllable duffing oscillator system [32]. This design allows us to tune the global phase space dynamics via a single parameter $\delta$ within the same system to avoid geometric confounds inherent when comparing distinct chaotic systems. Each input dimension initializes a three-dimensional dynamical system that evolves for time $T$ with $N$ discrete steps, producing trajectories $\phi_{\delta}(\mathbf{x},t)\in\mathbb{R}^{d\times N\times 3}$. The parameter $\delta$ controls the system’s phase space behavior:
@@
Architecture-Level Constraints. For tasks with inherently temporal inputs (Experiment 3), we leverage the SNN’s intrinsic dissipation of Leaky Integrate-and-Fire (LIF) neurons [19]. The membrane potential evolves as:
@@
where the leak parameter $\beta$ controls the effective integration window ($\tau_{\text{mem}}=-1/\ln(\beta)$), acting as the architectural counterpart to the encoding parameter $\delta$: high $\beta$ ($\approx 1.0$) preserves information across timesteps (weak constraint), while low $\beta$ ($\approx 0.1$) rapidly dissipates internal states (strong constraint). This setup aims to validate that our findings derive from the constraint principle itself, rather than properties specific to input encoding.
2.2 Experiment 1: Generalization in Cross-Encoding Classification
@@
We first investigate representational transfer [62, 16] as a direct proxy for learned invariance. The rationale is straightforward: if dissipative constraints lead to the abstraction of robust representations, a network trained on this specific dynamical regime should generalize to others, analogous to domain adaptation but operating across the spectrum of temporal dynamics.
@@
Specifically, we designed a cross-encoding evaluation protocol using the handwritten digits dataset (64 features, 10 classes) [2]. SNNs with LIF neurons are benchmarked against recurrent architectures and static baselines: Long Short-Term Memory networks (LSTMs) [26], Recurrent Neural Networks (RNNs), Multi-Layer Perceptrons (MLPs) processing only the final timestep (Last-T), and MLPs processing time-averaged inputs (Avg-Pool). All networks share an identical three-layer, fully-connected topology to isolate algorithmic differences. We sampled 12 dynamical regimes from $\delta\in[-1.5,10.0]$ and constructed a full $12\times 12$ generalization matrix, training each architecture on one regime ($\delta_{\text{train}}$) and evaluating across all others ($\delta_{\text{test}}$). Detailed experiment settings are provided in Appendix A.1.
@@
Figure 1: Cross-encoding generalization with hierarchical variance reduction. (a-e) 3D generalization landscapes (mean accuracy with deviation, $n=10$). (c) SNNs exhibit a robust generalization in the transition regime ($\delta\in[0,2]$) across the spectrum. This contrasts with the diagonal-only ridges seen in expansive regimes ($\delta<0$) and in baseline architectures (Last-T MLP, LSTM, RNN; b, d, e), and Avg-Pool MLP (a) shows partial, unstable generalization (see Appendix Tables 69). (f) Out-of-Distribution accuracy strongly anti-correlates with Layer 1 neural variability (CV; r=0.962r=-0.962). Error bars: $\pm 1$ SEM. (g) Hierarchical variance reduction (L1 $\to$ L3) emerges in SNN across regimes.
@@
Asymmetric Generalization Landscapes. Figure 1 presents generalization performance as a 3D topographic landscape. We define In-Distribution (ID) accuracy as the performance on the training encoding (the diagonal ridges where $\delta_{\text{train}}\approx\delta_{\text{test}}$), and Out-of-Distribution (OOD) accuracy as the performance across the remaining spectrum. While all architectures maintain high ID performance (ridge peaks), SNNs (Figure 1c) reveal a unique asymmetry: Networks trained in expansive regimes ($\delta_{\text{train}}<0$) form sharp, narrow ridges that achieve high ID accuracy but collapse immediately when $\delta_{\text{test}}>0$. This behavior is characteristic of memorization without transfer. In contrast, SNNs trained in the transition ($\delta_{\text{train}}\in[0,2]$) and dissipative ($\delta_{\text{train}}>2$) regimes form a robust "generalization plateau," maintaining high accuracy across the entire $\delta_{\text{test}}$ axis. Notably, the transition zone achieves the strongest overall generalization, effectively decoding inputs from all dynamical regimes.
This plateau is absent in recurrent controls (LSTMs, RNNs) and stateless MLPs (Last-T), which show OOD failure (collapsed ridges) when tested away from their training diagonal. The Avg-Pool MLP shows partial generalization but with significantly higher volatility (wider confidence bands in Figure 1a, see also Appendix Tables 59). This comparison reveals a key distinction: Avg-Pool achieves OOD robustness by collapsing input temporal variation before learning occurs, while SNNs induce invariance through the network learning process. The former provides a shortcut that discards information indiscriminately, which explains its higher performance volatility; the latter filters transient noise while retaining task-relevant features. Although LSTMs and RNNs preserve temporal autocorrelations, their poor performance suggests that generalization arises from alignment between signal dynamics and architecture, not temporal processing per se.
@@
Emergent Hierarchical Variance Reduction. How do SNNs translate these dynamical constraints into robust representations? We hypothesize that dissipative constraints drive the network to suppress noise while amplifying stable features. To test this, we measured neural response stability via the coefficient of variation ($CV=\sigma/\mu$) of mean firing rates across test encodings. Figure 1f reveals a mechanistic insight: response variability (CV) strongly anti-correlates with OOD accuracy across the dynamical spectrum (r=0.96r=-0.96; see Appendix Table 10).
@@
Furthermore, this stabilization emerges hierarchically across layers. As shown in Figure 1g, SNNs exhibit monotonic reduction with depth (L1 $\to$ L3) with the deepest layer showing the starkest contrast: when the training regime is under expansive dynamics, Layer 3 exhibits high fluctuating variance ($0.470\pm 0.288$, see Appendix Table 11), suggesting that the network memorizes transient trajectories of input signals. In contrast, dissipative constraints drive representation toward a tightly converged state ($0.307\pm 0.051$), indicating invariant feature formation (Appendix Table 11). This phenomenon is highly similar to the processing hierarchy observed in both standard deep networks [62] and the biological vision system (primate ventral stream), where raw, stimulus-specific volatility (akin to V1 responses) is progressively condensed into abstract representations (as in the inferotemporal cortex) [17, 1]. However, previous works primarily attribute such invariance to prior spatial architecture designs (e.g., pooling or strided convolutions [61]); our results demonstrate that hierarchical robustness can emerge from intrinsic temporal dissipation alone.
2.3 Experiment 2: Emergence of Structured Features
@@
The preceding findings naturally raise our next question: Does variance reduction spontaneously induce organized feature representations? Classical theories posit sparse coding as an explicit optimization objective for cortical receptive fields [43, 7]. We hypothesize instead that dissipative dynamics act as a fundamental constraint that naturally induces the invariant, structured features. In other words, representational structure is not the result of a designed objective but an emergent property of phase space compression.
To test this, we trained an SNN autoencoder to reconstruct static image patches from temporally encoded inputs. The architecture comprises a LIF bottleneck layer between a linear encoder ($W_{\text{enc}}$) and decoder ($W_{\text{dec}}$), where the decoder reconstructs from the temporal sum of spikes: $\mathbf{x}_{\text{recon}}=W_{\text{dec}}\sum_{t=1}^{T}S(t)$. Following established schemes [4, 22], we compared eight temporal encoding strategies on CIFAR-10 patches [33]: Baseline (static), Random (temporal jitter), Linear (rate coding), Poisson (stochastic spiking), and four dynamical regimes (Expansive, Critical, Transition, Dissipative).
The network minimized a composite loss balancing reconstruction fidelity and code sparsity (see Appendix A.2 for details). To evaluate learned representations, we measured the standard deviation of bottleneck receptive fields ($\sigma_{\text{RF}}$), which distinguishes structured filters from low-variance noise. Robustness was validated through: (1) varying the sparsity weight $\lambda$ (specifically including unconstrained $\lambda=0$); (2) scaling experiments to high-resolution STL-10 images [15]; and (3) testing dynamical universality across alternative systems (Lorenz [38] and Thomas attractors [53]). Detailed experiment protocols are provided in Appendix A.2.
@@
Structure Emergence under Dynamical Constraints Figure 2 provides a qualitative visualization of feature emergence. Transition regime ($\delta\approx 2.0$) spontaneously produces receptive fields with clear spatial organization: localized, oriented filters resembling the biological orientation maps. In contrast, baseline, random, and expansive conditions yield homogeneous, unstructured weight distributions resembling salt-and-pepper noise. Table 1 quantifies these distinct organizations across three metrics:
@@
  • Expansive dynamics ($\delta=-1.5$): Achieves excellent reconstruction ($0.0013$) but catastrophic sparsity ($1.72$), which means that dynamics amplify temporal variations, enabling memorization via high-rate firing without learning compressive features.
@@
  • Poisson encoding: Demonstrates the opposite extreme with near-perfect sparsity ($10^{-5}$) but minimal structure ($0.049$). This indicates that sparse firing alone is insufficient; temporal stability is required to induce structural differentiation.
@@
  • Transition dynamics ($\delta\approx 2.0$): Achieves optimal balance. Contractive phase space simultaneously reduces activation volume (sparsity $0.0009$), suppresses noise to enable clean filters, and amplifies stable features (structure $0.206$, $\mathbf{4.9\times}$ higher than alternatives).
@@
Further analysis confirms that even without sparsity penalty ($\lambda=0$, Table 12, Appendix A.2), the Transition regime still spontaneously induces structured representations ($\sigma_{\text{RF}}\approx 0.177$), whereas others collapse into noise. This structural emergence is robust: replication with Lorenz and Thomas attractors yields similar structured receptive fields exclusively within their dissipative regimes (Fig. 3, Appendix 3), confirming that this emergence is not specific to our oscillator system; a high-resolution parameter sweep reveals that this structural emergence persists as a stable ridge across the dynamical landscape (Fig. 4, Appendix A.2).
@@
Mechanism What specific physical property drives this structural emergence? Spectral analysis reveals that the transition regime ($\delta\approx 2$) occupies a unique computational niche defined by a "High-Entropy, Low-Frequency" signature (Fig. 5, Appendix 4). This signature is scale-invariant across timescales ($T_{\max}$) and sampling densities ($N$), ensuring that the network receives information-rich structure within a slowly-varying envelope. By restricting input trajectories to a low spectral centroid, transition dynamics function as an intrinsic spectral shaper that aligns input complexity with the spectral bias of neural networks [48], which also explains the asymmetric generalization observed in Experiment 1. These spectral properties confirm that feature organization arises solely from the dissipative dynamics itself, acting as a principle that induces structure more naturally than hand-crafted, explicit regularization.
@@
Functionally, the observed structure (Figure 2) mirrors the spatial antagonism found in biological vision [43, 20]. By suppressing high-frequency noise while preserving low-frequency structure, transition dynamics amplify spatiotemporal correlations. Such selective amplification biases the optimization process toward grouping co-varying features, yielding outcomes that are much like Hebbian association [23, 7], and it is driven by input temporal statistics rather than explicit plasticity rules. Conversely, high-frequency chaos in the expansive dynamics decorrelates inputs and prevents consolidation over time, while static baselines lack the necessary temporal continuity for structural binding. This suggests that dissipative constraints serve as a computational role beyond energy efficiency: by naturally filtering temporal correlations, they provide a mechanistic basis for structured connectivity [12]. More broadly, it provides a mechanistic account for biological sparse coding: under dissipative metabolic constraints, structured representations emerge not as a pre-programmed objective, but as a canonical solution to learning under temporal constraints [3].
@@
Beyond biological plausibility, this interpretation connects to recent work on diffusion dynamics [11], which identifies a critical gap between memorization and generative abstraction. Our transition regime positions learning on the generative side of this divide: expansive dynamics with high-frequency chaos allow the network to memorize transient noise for rapid error reduction (the "memorization phase"), whereas the transition regime imposes a temporal information bottleneck, which makes the system discard transient variations and rely on the underlying generative invariants (Table 4 and Fig. 5 in Appendix). This effectively implements Slow Feature Analysis [59] through dynamics alone, reframing the "performance valley" [13] not as a failure, but as the necessary cost of abstraction.